# ATPrompt: Textual Prompt Learning with Embedded Attributes

### v1

@@ 3.3 Attribute search
Table 1: Base-to-novel generalization experiments of five baselines with and without our ATPrompt on 11 recognition datasets. HM: Harmonic Mean. $\Delta$: HM improvement of ATPrompt over previous results. “ATPrompt” is abbreviated as “ATP”. Our method achieves consistent average performance improvement over different baselines.
Source Target Dataset
@@
Version Method Image Caltech Oxford Stanford Flowers Food FGVC SUN DTD Euro UCF Avg. $\Delta$
Net 101 Pets Cars 102 101 Aircraft 397 SAT 101
Shallow CoOp 71.51 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.92 46.39 66.55 63.88
@@
Table 2: Cross-dataset generalization experiments of three baselines with and without our ATPrompt on 11 datasets. Our method achieves consistent average performance improvements over three baseline methods.
We apply the alternating algorithm [11, 23] to solve this problem, fixing one set of variables and solving for the other set. Formally, we can alternate between solving these two subproblems:
@@
where $L_{train}$ and $L_{val}$ uses cross-entropy as the loss function. At the end of the search, an attribute can be selected from the ones with the highest weights.
Overall. The entire process comprises two steps. The first step is the attribute search stage. For downstream data, the attribute pool $\mathcal{V}$ is initially formed by LLMs. Subsequently, the attribute search method is conducted within this pool to identify the attribute $v$ that best suits our form. After the search is complete, the selected attributes are integrated into ATPrompt for targeted model training.
@@ 4.1 Settings
Base-to-Novel Generalization. Following [50, 49, 16, 24], we split the dataset into base and novel classes. The model is trained on the base class training set and evaluated on the test set. The attributes applied in this experiment are searched based on the base class. For datasets not including validation sets, such as ImageNet, we evenly divide the available 16-shot data, allocating half for training and the remaining half for attribute search.
Cross-dataset Experiments. Same as previous works [50, 49, 16], we first train a model on source dataset (ImageNet-1K) and then evaluate on out-of-distribution datasets to test generalization performance. The attributes adopted in this experiment are searched based on the source dataset.
@@
Attribute Search. Wxe select five independent attributes as the basis in the attribute pool. We do not particularly consider the order of attributes, as the following experiments verify that it will not significantly affect the training results. Therefore, for five attributes, we have 31 attribute combinations as candidates for search. We use ChatGPT-4o for attribute queries. Note that our proposed attribute search process does not take days like the traditional NAS method. Since there are very few parameters to be trained in this method, the search process only takes about 40 minutes on a V100 GPU for a dataset like Caltech101. In Tab. 3, we report a subset of the attribute bases queried from the LLM and the results obtained after searching. In the appendix, we show the specific weights learned for each candidate after the search process.
Implementation Details. We evaluate the effectiveness of our method on 11 recognition datasets, including ImageNet-1K [6], Caltech-101 [7], OxfordPets [30], StanfordCars [18], Flowers-102 [29], Food-101 [3], FGVCAircraft [27], SUN-397 [44], DTD [5], EuroSAT [8] and UCF-101 [36]. . The ViT-B/16 CLIP is selected as our default model. We report base and novel class accuracy and their harmonic mean (HM) averaged over 3 runs. Please refer to the Appendix for more experimental details.
@@ 4.2 Base-to-Novel Generalization
As demonstrated in Tab. 1, we evaluate the base-to-novel generalization performance of five baseline methods, both with and without integrating ATPrompt, across 11 recognition datasets. Notably, ATPrompt consistently enhances the average performance over all baseline methods.
Reasons for Limited Improvement in Some Conditions. (1) When we implement our method, we do not carefully adjust the hyperparameters of the baseline (e.g., learning rate, training epochs, loss hyperparameters); they remain consistent with the original method. (2) Recent advanced methods have integrated a variety of modules and regularization techniques (e.g., KG loss in KgCoOp, GPA, and SCL in PromptSRC). These additions may render the improvements introduced by our ATPrompt less perceptible or potentially redundant alongside existing enhancements. (3) Performance fluctuations due to limited training data.
@@ 4.3 Cross-dataset Evaluation
Tab. 2 shows the cross-dataset generalization results for three baseline methods. Our method demonstrates superior performance, with improvements of $1.38\%$, $0.85\%$ and $0.45\%$ for CoOp, CoCoOp and MaPLe respectively.
@@
4.4 Further Analysis
In this section, we conduct ablation experiments to evaluate the operations and hyperparameters of ATPrompt, thereby validating its effectiveness. By default, the experiments are conducted on the ImageNet. To minimize the influence of other components, we mainly adopt CoOp as the baseline method. Two attributes are used in our ATPrompt. For more experimental results, please refer to the Appendix.
Soft Prompt Length. In Fig. 5, we examine the optimal number of soft tokens for attribute and class tokens, respectively. By varying the length from 1 to 10, we can observe that introducing more soft tokens weakens the effect of embedded attribute tokens and reduces the generalization performance for unknown classes.
Class Token Position. Since our method introduces a new textual prompts format, the relative positioning of attribute tokens and class tokens requires careful consideration. In Tab. 4, we examine configurations where the class token is positioned in the middle or on either side of two attribute tokens. The results demonstrate that optimal performance is achieved when the class token is placed at the end, which aligns with the conclusions of CoOp.
@@ 4.4 Further Analysis
The results show that maintaining the attributes of both hard and soft tokens during the forward process results in optimal performance. Conversely, the actions of discarding and reintroducing attribute tokens diminish the model’s effectiveness. A possible explanation for this is that such operations disrupt the continuity of attribute representations across layers and amplify the disparity with existing tokens, thereby complicating the training of soft tokens.
Attribute Order. In this study, we do not specifically focus on the order of attributes in ATPrompt because varying the sequence usually does not result in semantic deviations in reality. For example, phrases like “a yellow round leaf” and “a round yellow leaf” convey the same meaning. Tab. 6 quantitatively assesses the impact of attribute order on prompt learning performance. From this table, we observe that despite variations in order, similar results are consistently produced, and the performance fluctuations across different orders remain within a reasonable range.
@@
Table 6: Comparison of different attribute orders on ImageNet. The order of attributes does not significantly affect the model, and performance fluctuations are within a reasonable range.
Comparison to Other Attributes. In Tab. 7, we explore the effectiveness of attributes derived through alternative methods, specifically by manually selecting class-irrelevant and common attributes. To highlight the comparison, we choose the fine-grained Food101 dataset for experiments.
@@
Common (shape, size) 88.48 86.87 87.67
(color, texture) 88.50 87.17 87.83
Irrelevant (plane, engines) 88.64 86.31 87.46
(football, sport) 88.60 86.18 87.37
(leather, silk) 88.56 86.82 87.68
Searched (flavor, preparation) 88.74 87.44 88.09
Table 7: Comparison of different attributes on Food101. The attributes obtained by our method achieve the best performance.
The results indicate that manually selected irrelevant attributes exhibit comparable performance during training; however, they perform poorly when applied to new categories. This suggests that incorrect attribute tokens cause the soft tokens to develop biased representations, thereby diminishing their zero-shot generalization ability.
@@ 5 Conclusion
In this work, we propose to utilize universal attributes as a bridge to enhance the association of images and unknown categories in prompt learning. Specifically, we introduce an attribute-embedded textual prompt learning method named ATPrompt. This approach expands the learning space of soft prompts from the original one-dimensional category level into a multi-dimensional attribute level by embedding fixed universal attribute tokens into the soft prompts. To identify the appropriate attributes, we develop an automated pipeline that learns to select suitable attributes from a candidate pool summarized by LLMs. Both shallow and deep versions of ATPrompt are introduced to ensure compatibility with various existing methods. Extensive experiments demonstrate the effectiveness of our approach.
Acknowledgement. This work was supported by the Young Scientists Fund of the National Natural Science Foundation of China (Grant No.62206134), the Fundamental Research Funds for the Central Universities 070-63233084 and the Tianjin Key Laboratory of Visual Computing and Intelligent Perception (VCIP). Computation is supported by the Supercomputing Center of Nankai University (NKSC). This work was supported by the National Science Fund of China under Grant Nos. 62361166670 and U24A20330. This work was also supported by DAMO Academy through DAMO Academy Research Internship Program.
@@ 6.1 Dataset
We evaluate the performance of our method on 15 recognition datasets. For generalization from base-to-novel classes and cross-dataset evaluation, we evaluate the performance of our method on 11 diverse recognition datasets. Specifically, these datasets include ImageNet-1K [6] and Caltech-101 [7] for generic object classification; OxfordPets [30], StanfordCars [18], Flowers102 [29], Food101 [3], and FGVCAircraft [27] for fine-grained classification, SUN-397 [44] for scene recognition, UCF-101 [36] for action recognition, DTD [5] for texture classification, and EuroSAT [8] for satellite imagery recognition. For domain generalization experiments, we use ImageNet-1K [6] as the source dataset and its four variants as target datasets including ImageNet-V2 [34], ImageNet-Sketch [40], ImageNet-A [10], and ImageNet-R [9].
@@ 6.2 Attribute Search
Inspired by DARTS [25], we utilize a differentiable search method to learn and select the appropriate attribute content and quantity for our proposed attribute-embedded form. During the search process, the batch size is set to 32, and the number of training epochs is set to 40. We use SGD as the optimizer for soft prompts $\theta$, as shown in Eqn. (12), with an initial learning rate of 0.002. Additionally, we employ Adam as the optimizer for the weight vector $\alpha$, as indicated in Eqn. (11), with an initial learning rate of 0.02. In our experiments, we configure the number of attribute bases to 5, resulting in the generation of 31 attribute combinations.
In Tab. 12, we present the five specific attribute bases obtained after querying the LLM, as well as the optimal attribute for training after searching. Tab. 13 displays the attributes and their corresponding weights from the final epoch of the search stage for the Caltech-101 dataset.
@@ 6.3 Base-to-Novel Generalization
Baseline Methods. We apply our ATPrompt to a wide range of textual-based prompt learning approaches, including CoOp [50], CoCoOp [49], KgCoOp [45], MaPLe [15], and PromptSRC [16].
Settings. Our code is implemented using PyTorch [31]. All experiments were conducted on a single NVIDIA V100 GPU. We adopted the standard data augmentation scheme as the baseline method, including random resized cropping and flipping. We employed stochastic gradient descent (SGD) as the optimizer. By default, we set the soft prompt lengths $a_{m}$, $b_{m}$, and $M$ in Eqn. (4) to be the same, as attribute and class tokens are considered equally important. The implementation details for each baseline method are presented as follows:
Figure 6: Architectural comparison between CoOp & KgCoOp and CoOp & KgCoOp + ATPrompt.
CoOp+ATPrompt: The experimental settings are consistent with the baseline method, with a batch size of 32 and an initial learning rate of 0.002. According to the original paper, the learnable prompt length $M$ for ResNet-50 CLIP is reported as 16; however, detailed results for the ViT-B/16 CLIP were not provided. In our setup, we configured the learnable attribute prompt lengths $a_{m}$ and $b_{m}$, as well as the learnable class prompt length $M$ to be 2. The baseline model is trained for 200 epochs using cosine decay. In our approach, we train the model for 100 epochs while maintaining the same decay schedule. Fig. 6 illustrates the architectural comparison between the original CoOp and the CoOp+ATPrompt.
KgCoOp+ATPrompt: In accordance with the baseline method, we use a batch size of 128 and an initial learning rate of 0.002. The loss hyperparameter $\lambda$, as reported in the paper, is set at 8. However, given that our ATPrompt already exhibits a regularization effect, we reduce this hyperparameter to 2. We adhere to the same training schedule as the baseline method. A detailed architectural comparison between KgCoOp and KgCoOp+ATPrompt is presented in Fig. 6.
CoCoOp+ATPrompt: In line with the baseline method, we employ a batch size of 1 and an initial learning rate of 0.002. The original paper specifies a learnable class prompt length of 4. In contrast, our method sets the lengths of both the learnable attribute soft prompts, $a_{m}$ and $b_{m}$, as well as the learnable class soft prompt, $M$, to 2. We follow the same training schedule as the baseline method, which involves training the model over 10 epochs using cosine decay.
In the original paper, CoCoOp introduces a meta network that processes image features to generate meta tokens, which serve as offsets for all soft prompt tokens. In our CoCoOp+ATPrompt, we maintain both the meta network and the meta tokens; however, the meta tokens are exclusively used as offsets for class soft tokens $[T_{1}],...,[T_{M}]$, as illustrated in Fig. 7.
@@
Version Method Source Target Dataset
ImageNet -V2 -S -A -R Avg. $\Delta$
Shallow CoOp 71.51 64.20 47.99 49.71 75.21 59.28
+ATPrompt 71.67 64.43 49.13 50.91 76.24 60.18 (+0.90)
Shallow CoCoOp 71.02 64.07 48.75 50.63 76.18 59.91
+ATPrompt 71.27 64.66 49.15 51.44 76.33 60.40 (+0.49)
Deep MaPLe 70.72 64.07 49.15 50.90 76.98 60.27
+ATPrompt 70.69 64.40 49.10 51.77 77.11 60.60 (+0.33)
Table 8: Domain generalization experiments of three baselines with and without our ATPrompt on 4 datasets. Our method achieves consistent average performance improvement over three baseline methods.
MaPLe+ATPrompt: Following the baseline method, we use a batch size of 4 and an initial learning rate of 0.0035. While the original paper sets the length of the learnable prompt to 2, our approach involves setting the lengths of both the learnable attribute soft prompts, $a_{m}$ and $b_{m}$, as well as the learnable class soft prompt, $M$, to 4. Our training schedule mirrors that of the baseline method.
According to the original paper, all textual soft tokens are input into the projection layer to acquire visual tokens, which are subsequently inserted into the deeper layers of the image encoder. However, in MaPLe+ATPrompt, we only input class soft tokens into the projection layer and do not modify the attribute soft and hard tokens. A detailed architectural comparison between MaPLe and MaPLe+ATPrompt is illustrated in Fig. 8.
@@
PromptSRC+ATPrompt: Similar to the baseline method, our approach uses a batch size of 4, an initial learning rate of 0.0025, a text loss weight of 25, and an image loss weight of 10. We adhere to the same training schedule as the baseline method. In the original study, the soft prompt length is specified as 4. In our approach, we set the lengths of both the learnable attribute prompts $a_{m}$ and $b_{m}$, as well as the learnable class soft prompt $M$, to 4.
In our study, we exclusively remove and later reintroduce deep class soft tokens $[T_{1}]_{i},...,[T_{M}]_{i}$, as explained in our paper. Meanwhile, attribute-related soft tokens $[T_{a_{1}}]_{i},...,[T_{b_{m}}]_{i}$, along with hard tokens $[A]_{i}$ and $[B]_{i}$, are retained throughout the forward process. A comprehensive architectural comparison between PromptSRC and PromptSRC+ATPrompt is presented in Fig. 9.
Figure 9: Architectural comparison between PromptSRC and PromptSRC+ATPrompt. In PromptSRC+ATPrompt, we only drop and reintroduce class soft tokens in the deep layers. The attribute soft and hard tokens are retained throughout the forward process.

### v2

@@ 3.3 Attribute search
Table 1: Base-to-novel generalization experiments of five baselines with and without our ATPrompt on 11 datasets. Our method achieves consistent average performance improvement over different baselines.
Method Source Target Dataset Average
@@
Image Caltech Oxford Stanford Flowers Food101 FGVC SUN397 DTD Euro UCF101
Net 101 Pets Cars 102 Aircraft SAT
CoOp 71.51 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.92 46.39 66.55 63.88
@@
Table 2: Cross-dataset generalization experiments of three baselines with and without our ATPrompt on 11 datasets. ATPrompt achieves consistent average performance improvement on target datasets.
@@
where $L_{train}$ and $L_{val}$ both use the cross-entropy loss function. After the search, the attribute combination with the highest weight ($\alpha_{i}$) is selected.
Cost Analysis. Unlike traditional Neural Architecture Search (NAS) methods [29, 8, 44], which search computationally expensive network-level parameters, our approach focuses on a lightweight token-level search space. This design makes our method significantly more efficient than previous approaches. In practice, our search converges in approximately 5 epochs, oftern requiring less than 5 minutes on a single A800 GPU. This highlights the practical advantages of our method over traditional parameter architecture search. In addition, search efficiency can be improved by curating a smaller set of base attributes, which narrows the search space.
@@ 4.1 Settings
Base-to-Novel Generalization. Following [59, 58, 19, 27], we split the dataset into base and novel classes. The model is trained on the base class training set and evaluated on the test set. The attributes applied in this experiment are searched based on the base class. For datasets without a dedicated validation set, like ImageNet, we split the 16-shot labeled data, using half for training and the other half for attribute search.
Cross-dataset Experiments. Consistent with previous works [59, 58, 19], we first train a model on the ImageNet-1K source dataset and then evaluate its generalization performance on several out-of-distribution datasets. The attributes used are obtained from the source dataset.
@@
Food101 flavor, texture, origin, ingredients, preparation (flavor, preparation)
@@
Attribute Search. We select five independent attributes as the basis in the attribute pool. This results in 31 candidate attribute combinations for the search process. We use ChatGPT-4o for attribute queries. In Tab. 3 reports a subset of the queried attribute bases and the final combination selected by our search. The Appendix provides further details on our search process.
Implementation Details. We evaluate the model performance on 15 popular recognition datasets. We report base and novel class accuracy and their harmonic mean (HM) averaged over 3 runs. The details of each dataset are attached in the Appendix.
@@ 4.2 Base-to-Novel Generalization
As demonstrated in Tab. 1, we evaluate the base-to-novel generalization performance of five baseline methods, both with and without integrating ATPrompt, across 11 diverse recognition datasets. Notably, ATPrompt consistently improves the average performance of all baseline methods.
Reasons for Limited Improvement in Some Conditions. (1) Learnable text prompts are the central component in earlier works, and optimizing them through ATPrompt can obviously improve the performance. (2) Recent studies have expanded beyond learnable text prompts by introducing additional learnable modules. Since our work only involves the optimization of the learnable text prompt part, the improvement have become less obvious.
@@ 4.3 Cross-dataset Evaluation
Tab. 2 shows the cross-dataset generalization results for three baseline methods. Our method demonstrates superior performance, with improvements of $1.38\%$, $0.85\%$ and $0.45\%$ for CoOp, CoCoOp and MaPLe, respectively.
4.4 Domain Generalization
Tab. 4 shows the domain generalization results for three baseline methods. The results show that our method improves CoOp, CoCoOp and MaPLe methods by $0.90\%$, $0.49\%$ and $0.33\%$ respectively.
Method Source Target Dataset Average
ImageNet -V2 -S -A -R
CoOp 71.51 64.20 47.99 49.71 75.21 59.28
+ATPrompt 71.67 64.43 49.13 50.91 76.24 60.18 (+0.90)
CoCoOp 71.02 64.07 48.75 50.63 76.18 59.91
+ATPrompt 71.27 64.66 49.15 51.44 76.33 60.40 (+0.49)
MaPLe 70.72 64.07 49.15 50.90 76.98 60.27
+ATPrompt 70.69 64.40 49.10 51.77 77.11 60.60 (+0.33)
Table 4: Domain generalization experiments of three baselines with and without our ATPrompt on four datasets. The integration of ATPrompt resulted in better generalization performance.
4.5 Further Analysis
By default, the experiments are conducted on the ImageNet. To minimize the influence of other components, we mainly adopt CoOp as the baseline method. Two attributes (color and shape) are used in our ATPrompt. For more experimental results, please refer to the Appendix.
@@ 4.5 Further Analysis
Soft Prompt Length. In Fig. 5, we examine the optimal soft token length for both attribute and class tokens. By varying the length from 1 to 10, we observe that longer prompts dilute the guiding influence of the attribute tokens, thereby reducing generalization to novel classes.
Class Token Position. The relative positioning of attribute tokens and class tokens in our method requires careful consideration. In Tab. 5, we examine configurations where the class token is positioned in the middle or on either side of two attribute tokens. The results demonstrate that optimal performance is achieved when the class token is placed at the end, which aligns with the findings of CoOp.
@@
The results show that maintaining the attributes of both hard and soft tokens during the forward process results in optimal performance. Conversely, dropping and reintroducing either hard or soft attribute tokens harms performance, likely because it disrupts the continuity of attribute representations across layers and complicates optimization.
@@
Table 7: Comparison of different orders on ImageNet. The order of attributes does not significantly affect the model, and performance fluctuations are within a reasonable range.
Attribute Order. In this study, we do not specifically focus on the order of attributes because varying the sequence usually does not result in semantic deviations in reality. Tab. 7 quantitatively assesses the impact of attribute order on prompt learning performance. From this table, we observe that despite variations in order, similar results are consistently produced, and the performance fluctuations across different orders remain within a reasonable range.
Comparison to Other Attributes. In Tab. 8, we explore the effectiveness of attributes derived through alternative methods, specifically by manually selecting class-irrelevant and common attributes. The results indicate that manually selected irrelevant attributes exhibit comparable performance during training; however, they perform poorly when applied to new categories. This suggests that incorrect attribute tokens cause the soft tokens to develop biased representations, thereby diminishing their zero-shot generalization ability.
@@
Common (shape, size) 82.83 67.13 74.16
(color, texture) 82.73 67.56 74.38
Irrelevant (plane, engines) 82.81 66.22 73.59
(football, sport) 82.77 67.14 74.14
Searched - 82.68 68.04 74.65
Table 8: Comparison of average performance for various attribute configurations on 11 datasets. The attributes obtained by our method achieve the best performance.
@@ 5 Conclusion
In this work, we introduce ATPrompt, an attribute-anchored textual prompt learning method that uses universal attributes as a bridge to improve generalization from seen to unseen categories. Our approach expands the learning space of soft prompts from a one-dimensional, category-centric structure to a multi-dimensional attribute space by anchoring fixed attribute tokens with the prompt. To ensure the selection of optimal attributes, we propose an automated pipeline designed to identify the most suitable candidates for any given downstream task. ATPrompt is designed with both shallow and deep architectural variants, rendering it broadly compatible with existing prompt learning methods. Extensive experiments validate the effectiveness of our approach. We believe this work offers a new direction for research into the fundamental structure of learnable prompts in prompt learning area.
Limitations and future works. This work is a preliminary study of the basic prompt form, which has not been able to achieve comparable performance to regularization-based methods when working alone. Furthermore, the current selection of attribute anchors relies on manual experimentation, and automatically discovering the optimal anchor positions through a learning-based approach remains a promising direction for future research.
Acknowledgments. This work was supported by the National Natural Science Foundation of China under Grant Nos. 62361166670 and U24A20330. This research was also supported by the Young Scientists Fund of the National Natural Science Foundation of China (Grant No.62206134), the Fundamental Research Funds for the Central Universities 070-63253222 and the Tianjin Key Laboratory of Visual Computing and Intelligent Perception (VCIP). This work was also supported by Shenzhen Science and Technology Program (JCYJ20240813114237048). Computation is supported by the Supercomputing Center of Nankai University (NKSC).
@@ 6.1 Dataset
We evaluate the performance of our method on 15 recognition datasets. For generalization from base-to-novel classes and cross-dataset evaluation, we evaluate the performance of our method on 11 diverse recognition datasets. Specifically, these datasets include ImageNet-1K [6] and Caltech-101 [9] for generic object classification; OxfordPets [35], StanfordCars [21], Flowers-102 [34], Food-101 [3], and FGVCAircraft [32] for fine-grained classification, SUN-397 [50] for scene recognition, UCF-101 [41] for action recognition, DTD [5] for texture classification, and EuroSAT [10] for satellite imagery recognition. For domain generalization experiments, we use ImageNet-1K [6] as the source dataset and its four variants as target datasets including ImageNet-V2 [39], ImageNet-Sketch [46], ImageNet-A [12], and ImageNet-R [11].
@@ 6.2 Attribute Search
Inspired by DARTS [30], we employ a differentiable search method to identify the optimal content and quantity of attributes for our proposed attribute-anchored form. The search process is conducted for 10 epochs with a batch size of 32. We use SGD to optimize the soft prompts $\theta$ with an initial learning rate of $0.002$. and Adam to optimize the weight vector $\alpha$ with an initial learning rate of $0.02$. In our experiments, we use 5 attribute bases, which generate 31 (i.e., $C_{5}^{1}+C_{5}^{2}+C_{5}^{3}+C_{5}^{4}+C_{5}^{5}$) candidate combinations for the search process.
Tab. 11 presents the five attribute bases generated by the LLM, alongside the optimal attribute combination identified after the search. Furthermore, Tab. 13 displays the final weights of all candidate combinations from the search stage on the Caltech-101 dataset.
@@ 6.3 Base-to-Novel Generalization
Baseline Methods. To evaluate ATPrompt, we integrate it with several leading textual-based prompt learning approaches, including CoOp [59], CoCoOp [58], MaPLe [18], DePT [55] and PromptKD [27]. The experimental settings are detailed below.
Settings. Our framework is implemented in PyTorch [36] and all experiments were conducted on a single NVIDIA A800 GPU. Following the baseline methods, we use a standard data augmentation scheme of random resized cropping and flipping. We employe Stochastic Gradient Descent (SGD) as the optimizer. By default, the soft token lengths for attribute and class tokens are set to be identical, as attribute and class tokens are considered equally important. The specific implementation details for each baseline method are presented as follows:
Figure 6: Architectural comparison between CoOp and CoOp+ATPrompt.
CoOp+ATPrompt: Following the baseline, we use a batch size of 32 and an initial learning rate of 0.002. The original paper reports a learnable prompt length of $M=16$ for ResNet-50 but does not specify a length for ViT-B/16. In our setup, we set the sofo token length for both the attribute and class tokens to 2. While the baseline model is trained for 200 epochs, we reduce the training to 100 epochs while maintaining the same cosine decay schedule. Figure 6 illustrates the architectural differences between the original CoOp and CoOp+ATPrompt.
CoCoOp+ATPrompt: We adhere to the baseline’s settings with a batch size of 1 and an initial learning rate of 0.002. Whereas the original paper specifies a soft class token length of 4, we set the length of our learnable tokens for the attribute and class token to 2. We adopt the same training schedule as the baseline: 10 epochs with cosine decay.
CoCoOp’s original design uses a meta-network to generate offsets for all soft prompt tokens. We retain this meta-network but modify its application: the meta tokens now serve as offsets exclusively for the class soft tokens, $[T_{1}],...,[T_{M}]$, as shown in Fig. 7.
@@
MaPLe+ATPrompt: We adhere to the baseline hyperparameters, utilizing a batch size of 4 and an initial learning rate of 0.0035. We diverge from the original prompt configuration; whereas the baseline sets the learnable prompt length to 2, our method sets the soft token lengths of both the attribute and the class token to 4. The training schedule remains consistent with the baseline.
The primary architectural modification in MaPLe + ATPrompt concerns the projection mechanism. The original MaPLe framework inputs all textual soft tokens into a projection layer to generate corresponding visual tokens, which are then fused into the image encoder. Our approach, however, selectively inputs only the class soft tokens into this projection layer, while the attribute tokens are preserved without modification. This architectural difference is visualized in Fig. 8.
DePT+ATPrompt: We adopt the baseline’s training configuration: a batch size of 32, a learning rate of 0.0035, a balance weight of $\lambda$=0.7, and a duration of 10 epochs. Our primary configuration for DePT+ATPrompt uses a learnable token length of 4. For datasets with lower complexity, namely Caltech-101, OxfordPets, and StanfordCars, we adjust these parameters, setting the soft token length to 2 and the balance weight to 0.6. The architectural differences between the DePT and DePT+ATPrompt models are detailed in Fig. 9.
@@
Figure 9: Architectural comparison between DePT and DePT+ATPrompt.